OpenStudy (anonymous):

The diameter of a circle is 16cm By what amount must the radius be decreased in order to decrease the area by 48pi square centimeters. (i have to solve it using applications of equations)

OpenStudy (anonymous):

area = pi*d/4

OpenStudy (anonymous):

area = pi*(d^2)/4

OpenStudy (anonymous):

sry...

OpenStudy (anonymous):

how did you get this?

OpenStudy (anonymous):

that is just the equation for the area of a circle..

OpenStudy (anonymous):

so to decrease the area...

OpenStudy (anonymous):

i understand that, but how do i get waht number the radius decreases by?

OpenStudy (anonymous):

two equations

OpenStudy (anonymous):

area = pi*(16^2)/4 area -48pi = pi*(16^2)/4

OpenStudy (anonymous):

pi*(256/4)-48pi

OpenStudy (anonymous):

area -48pi = 64pi - 48 pi

OpenStudy (anonymous):

area - 48pi = 16pi

OpenStudy (anonymous):

16pi is your answer

OpenStudy (anonymous):

ugh im so confused :( i wish my teacher would acctually teach me stuff

OpenStudy (anonymous):

ok.. so lets backup

OpenStudy (anonymous):

grab a beer.. lets go...

OpenStudy (anonymous):

start with the equation of a circle

OpenStudy (anonymous):

area = pi (d^2/4)

OpenStudy (anonymous):

X 2 + Y 2=r2?

OpenStudy (anonymous):

now, both side are in "area" units...

OpenStudy (anonymous):

yes, that is the equation of a circle.. i meant the equation for the area fo a circle..

OpenStudy (anonymous):

area = pi (d^2/4)

OpenStudy (anonymous):

oh A= pi(r2)

OpenStudy (anonymous):

start with that guy

OpenStudy (anonymous):

yeah we can use that one too

OpenStudy (anonymous):

area = pi * r^2

OpenStudy (anonymous):

subtract from both side the area you want to change

OpenStudy (anonymous):

so A= pi(8^2)

OpenStudy (anonymous):

area -48pi = pi (8^2)-48pi

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh i get since we know the area we put it out front

OpenStudy (anonymous):

kinda.. we know what area we want to subtract... to we can put that on both sides of the equation without changing anything

OpenStudy (anonymous):

so break down that equation...

OpenStudy (anonymous):

area -48pi = 64pi -48pi

OpenStudy (anonymous):

then add thos together to get 16pi?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

then add 48 to both sides?

OpenStudy (anonymous):

now what radius gives you 16pi

OpenStudy (anonymous):

4

OpenStudy (anonymous):

pi r^2

OpenStudy (anonymous):

sqrt of 16 is 4

OpenStudy (anonymous):

easy peasy

OpenStudy (anonymous):

oh okay thanks i get it

OpenStudy (anonymous):

although im not sure how to put that in the way she wants us to write it

OpenStudy (anonymous):

that was always my problem too... i got the answers... just the work was never there..